OpenStudy (anonymous):

For the isomerization reaction: butane ⇌ isobutane Kp equals 25 at 500°C. If the initial pressures of butane and isobutane are 5.0. atm and 0.0 atm, respectively, what are the pressures of the two gases at equilibrium?

OpenStudy (mos1635):

butane ⇌ isobutane Po 0 -P +P Po-P P Kp=25 P/Po-P=25 P=25*(Po-P) P=(25/26)*Po P=125/26 atm so..........